# Ineq-Comp: Benchmarking Human-Intuitive Compositional Reasoning in Automated Theorem Proving on Inequalities

### v1

@@ Abstract
LLM-based formal proof assistants (e.g., in Lean) hold great promise for automating mathematical discovery. But beyond syntactic correctness, do these systems truly understand mathematical structure as humans do? We investigate this question through the lens of mathematical inequalities—a fundamental tool across many domains. While modern provers can solve basic inequalities, we probe their ability to handle human-intuitive compositionality. We introduce Ineq-Comp, a benchmark built from elementary inequalities through systematic transformations, including variable duplication, algebraic rewriting, and multi-step composition. Although these problems remain easy for humans, we find that most provers—including Goedel, STP, and Kimina-7B—struggle significantly. DeepSeek-Prover-V2-7B shows relative robustness —possibly because it is trained to decompose the problems into sub-problems— but still suffers a 20% performance drop (pass@32). Strikingly, performance remains poor for all models even when formal proofs of the constituent parts are provided in context, revealing that the source of weakness is indeed in compositional reasoning. Our results expose a persisting gap between the generalization behavior of current AI provers and human mathematical intuition.
@@ 1 Introduction
Parallel to progress in natural language reasoning, formal theorem proving with LLMs has developed rapidly in the past years. In particular, AlphaProof [^1] https://deepmind.google/discover/blog/ai-solves-imo-problems-at-silver-medal-level/ and AlphaGeometry (Trinh et al., 2024) showed that a model could achieve silver-medal performance in the International Math Olympiad by finding verifiable proofs. Open-source research efforts also show significant progress through both whole-proof generation and tree-search approaches (Xin et al., 2024; Lin et al., 2025; Dong and Ma, 2025; Wu et al., 2024; Li et al., 2024; Xin et al., 2025; Wang et al., 2025), marking rapid advancements in formal theorem proving.
@@
    We present Ineq-Comp, a benchmark for evaluating the compositional reasoning abilities of formal theorem provers on inequalities. It is created using 75 seed problems: Olympiad questions solved using basic tools such as AM-GM, Cauchy, and Jensen inequalities, each accompanied by a verified Lean 4 proof. These serve both as correctness guarantees and as valuable human-written formal data. Starting from these seed questions, we apply simple algebraic and compositional transformations (Figure 1, Left), generating 150 additional problems across two types of transformations (Section 2.1). Ineq-Comp is designed to be fully extensible: further composed inequalities can be automatically generated using rule-based transformations on any seed set (Section 2.2). We also include 50 real-world inequality problems to enhance diversity (Section 2.3).
@@
    We benchmark a range of LLM-based provers—from general-purpose models like R1-Distill-Qwen to tree-search-based systems—on Ineq-Comp. All models, except DeepSeek-Prover-V2-7B (Ren et al., 2025), suffer from significant performance drops even on the simplest transformations (Section 3). Notably, DeepSeek-Prover-V2-7B also shows over 20% drop when tested under a limited computational budget (pass@32). These failures persist even when the model is provided with the formal proof of the seed problem in-context (Section 4.1).
@@
    Through detailed output analysis, we identify two key limitations: (1) models rarely attempt to apply core inequalities such as AM-GM or Cauchy within their formal proofs, even when they refer to these techniques in natural language comments; and (2) most models rely heavily on a narrow set of tactics—particularly the sum-of-squares approach using nlinarith—indicating limited diversity in their proof strategies when it comes to inequalities (Section 3).
@@
    We test whether simple compositions in Ineq-Comp can be learned by fine-tuning synthetically generated problems (Section 4.2). While fine-tuning improves compositional performance on the training seed set, it fails to generalize to new seed problems. This suggests that compositional generalization in formal theorem proving is not easily learnable through data augmentation alone and may require fundamental advances in model architecture or training.
@@ 2 Dataset Curation
    Ineq-Simp (Section 2.1): 150 problems created by applying human-intuitive transformations to the 75 seed problems. These include algebraic rewrites and compositional constructions (e.g., duplicating variables or multiplying inequalities), as illustrated in Figure 1 (Left). Problems are categorized by the technique used in the seed (AM-GM, Cauchy, or Miscellaneous).
@@
    Ineq-Mix (Section 2.2): A rule-based, automated framework that generates new composed inequalities from any given seed set. This enables scalable extension of Ineq-Simp while preserving formal correctness.
@@
    Ineq-Real (Section 2.3): A set of 50 real-world inequality problems collected from math contests and educational resources to enhance benchmark diversity and assess generalization beyond synthetic construction.
@@ 2.2 Ineq-Mix: A fully expandable benchmark for composition
We introduce Ineq-Mix, an automated framework that generates a potentially unbounded set of valid inequality problems by applying predefined transformation rules to seed problems. These rules fall into three categories: (1) Compositional transformations (e.g., combining inequalities via addition, multiplication, max/min); (2) Variable-level transformations (e.g., replacing variables with algebraic expressions); (3) Problem-level transformations (e.g., applying a monotonic function like $\exp$ or $\log$ to both sides). All generated problems are guaranteed to be mathematically valid and formally solvable. Unlike the simpler variants in Section 2.1, Ineq-Mix produces more complex compositions that often require nontrivial decomposition and planning. For example, proving $f_{1}f_{2}\geq g_{1}g_{2}$ from $f_{1}\geq g_{1}$ and $f_{2}\geq g_{2}$ may require identifying intermediate inequalities. We generate 100 evaluation problems using only compositional rules (Table 3), which already challenge state-of-the-art provers. As models improve, Ineq-Mix can scale in complexity by combining multiple transformation types.
To assess whether this difficulty is inherent to formal reasoning, we test OpenAI’s O3 model, which operates in natural language, on problems involving one transformation from each category. O3 reliably decomposes and solves these using intuitive reasoning, with failures typically due to mistakes on individual subproblems. This contrast highlights a key gap: informal models like O3 handle compositionality with ease, while formal provers struggle even on problems built from elementary components. This suggests the bottleneck lies in current formal reasoning systems, not the mathematical content. See Appendix B for a full list of transformations used in Ineq-Mix.
@@ 3.1 Experiment setup
We evaluate three categories of models and methods: (1) General-purpose language models, (2) Whole-proof generation models, and (3) Tree-search methods. For general-purpose models, we test DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025) in two settings: with and without explicit thinking steps. Without the chat-style prompt, the model directly completes the Lean 4 proof code. In contrast, when provided with a chat template and generation prompt, the model first performs a “thinking process” in natural language before outputting the full Lean script. For whole-proof generation, we include DeepSeek-Prover-V1.5-RL (Xin et al., 2024), Goedel-Prover-SFT (Lin et al., 2025), STP (Dong and Ma, 2025), Kimina-7B (Wang et al., 2025), and the latest state-of-the-art model DeepSeek-Prover-V2-7B (Ren et al., 2025), which achieves top performance on the MiniF2F benchmark. For tree-search-based approaches, we evaluate DeepSeek-Prover-V1.5-RL+RMaxTS (Xin et al., 2024) and InternLM-2.5-Step-Prover+BF (Wu et al., 2024). While recent models such as HunyuanProver (Li et al., 2024) and BFS-Prover (Xin et al., 2025) show stronger performance on MiniF2F, we exclude them due to the lack of publicly available code.
@@
We evaluate all methods using the standard pass@$N$ accuracy, where a problem is considered solved if at least one of $N$ generated proofs is correct. For tree-search methods, budget denotes the total number of interactions with the Lean compiler. For additional details, see Appendix D, and refer to Appendix E for the prompt templates used in evaluation.
@@ 3.2 Main findings and discussion
The experiment results on Ineq-Simp: categorized by the natural of seed problems Ineq-AMGM, Ineq-Cauchy, and Ineq-MISC are summarized in Table 1. The performance of different models on Ineq-Mix (randomly generated) and Ineq-Real is shown in Figure 2.
@@
Our experiments reveal a clear failure of compositional generalization across most current theorem provers, even on problems derived from simple seed inequalities that remain easy for humans. As shown in Table 1, top models such as Goedel-Prover-SFT (Lin et al., 2025), STP (Dong and Ma, 2025), and Kimina-Prover (Wang et al., 2025) show large drops in accuracy from seed problems to their compositional variants: Type I (duplication-based) and Type II (algebraic transformation).
For instance, Kimina-Prover-Preview-Distill-7B achieves 80% on AM-GM seed problems (budget 3200), but drops to 44% on Type I and 64% on Type II. The drop is even more dramatic for Goedel-Prover and DeepSeek-Prover-V1.5, which struggle to exceed 5% accuracy on Ineq-AMGM Type I problems—even with a large number of attempts. The only partial exception is DeepSeek-Prover-V2-7B, which reduces the seed-to-Type I gap to 13% and seed-to-Type II to 18%. Its drop becomes more pronounced at smaller budgets (e.g., pass@32), confirming that the challenge is not just probabilistic but fundamentally hard (Figure 1, right). This is especially striking given that DeepSeek-Prover-V2 is explicitly trained using a divide-and-conquer strategy (Ren et al., 2025), which should, in principle, help with decomposing Type I problems. Its performance gap, particularly under low-pass settings, highlights the intrinsic difficulty of compositional reasoning in formal reasoning.
@@
Method Budget Ineq-AMGM Ineq-Cauchy Ineq-MISC Seed Type I Type II Seed Type I Type II Seed Type I Type II General Purpose Models DeepSeek-R1-Distill-Qwen-32B (w/o thinking)(Guo et al., 2025) 32 48.21.9 3.53.3 16.23.0 28.03.3 17.03.2 15.03.0 41.43.7 13.44.5 15.44.4 64 49.01.7 6.54.1 18.42.4 30.63.2 19.52.8 16.82.7 44.53.2 17.74.0 20.24.8 128 49.92.0 10.64.2 20.02.5 32.62.9 21.83.2 19.02.6 47.43.1 21.13.7 25.44.2 3200 52.0 40.0 36.0 44.0 32.0 28.0 52.0 36.0 36.0 DeepSeek-R1-Distill-Qwen-32B (w thinking)(Guo et al., 2025) 32 48.81.6 10.93.8 21.13.1 42.92.5 27.03.4 18.42.4 50.52.3 18.94.6 22.04.0 64 49.51.9 14.54.4 23.03.4 44.52.4 30.32.9 20.62.3 51.90.6 23.74.9 26.23.1 128 50.92.1 19.24.1 26.14.3 46.22.3 32.62.7 22.12.0 52.00.0 28.03.9 29.42.7 3200 60.0 44.0 44.0 56.0 40.0 24.0 52.0 36.0 40.0 Whole-Proof Generation Methods DeepSeek-Prover-V1.5-RL (Xin et al., 2024) 32 48.13.0 0.00.4 8.21.5 14.93.2 2.91.8 4.41.4 40.22.8 12.41.1 12.22.5 64 50.62.9 0.10.6 9.01.7 17.02.7 3.71.1 5.01.9 42.12.3 12.71.7 13.82.9 128 52.22.1 0.20.8 9.82.0 18.72.7 4.00.0 6.12.3 43.21.6 13.32.2 16.22.9 3200 60.0 4.0 24.0 24.0 4.0 12.0 44.0 20.0 28.0 Goedel-Prover-SFT (Lin et al., 2025) 32 48.62.9 0.41.2 14.03.2 34.82.5 12.43.5 21.53.4 47.01.7 14.43.1 24.61.9 64 50.62.6 0.81.6 16.62.8 36.21.9 15.83.4 24.62.9 47.80.9 16.62.5 25.51.9 128 52.21.4 1.31.9 18.62.2 37.11.8 19.42.9 26.91.8 48.00.0 17.92.6 26.42.5 3200 60.0 4.0 24.0 40.0 32.0 28.0 48.0 24.0 36.0 STP (w/o miniF2F valid) (Dong and Ma, 2025) 32 59.11.9 14.34.4 23.24.5 35.22.4 14.62.7 16.02.6 55.61.3 12.65.0 27.63.6 64 60.10.6 18.54.1 28.24.6 36.82.4 16.72.8 17.32.7 56.00.0 17.84.9 31.04.1 128 60.31.1 24.34.1 33.03.6 37.92.6 18.43.0 18.93.3 56.00.0 24.04.4 33.94.1 3200 64.0 44.0 40.0 44.0 24.0 28.0 56.0 36.0 40.0 Kimina-Prover-Preview-Distill-7B (Wang et al., 2025) 32 59.44.1 11.75.4 45.23.7 46.94.5 27.02.6 27.73.3 44.21.3 18.13.9 35.82.0 64 64.14.6 19.45.9 48.62.4 52.74.3 28.82.5 30.22.8 44.61.4 22.32.9 36.82.0 128 69.44.2 28.25.4 50.62.2 57.63.6 30.43.0 32.01.6 45.11.8 25.62.5 37.62.5 3200 80.0 44.0 64.0 68.0 52.0 36.0 52.0 32.0 44.0 DeepSeek-Prover-V2-7B (Ren et al., 2025) 32 75.04.4 58.64.0 52.54.5 64.64.1 33.02.3 35.02.3 59.12.9 49.33.4 38.84.4 64 80.75.3 62.14.5 57.44.0 68.33.1 34.72.7 36.62.3 61.72.5 51.62.9 43.74.2 128 85.85.4 65.95.3 61.43.7 71.02.0 36.33.6 37.92.6 64.01.6 53.33.1 49.94.3 3200 96.0 84.0 76.0 76.0 52.0 48.0 68.0 64.0 64.0 Tree Search Methods DeepSeek-Prover-V1.5-RL +RMaxTS(Xin et al., 2024) 1$\times$3200 60.00.0 3.01.7 22.02.0 24.00.0 8.02.8 13.03.3 44.00.0 14.03.5 29.01.7 2$\times$3200 60.00.0 6.02.0 26.02.0 24.00.0 10.02.0 16.00.0 44.00.0 16.04.0 32.00.0 4$\times$3200 60.0 8.0 28.0 24.0 12.0 20.0 44.0 20.0 36.0 InternLM2.5-StepProver+BF (Wu et al., 2024) 1$\times$32$\times$600 30.83.1 0.00.0 2.53.1 12.00.0 0.00.0 1.21.9 34.02.0 2.22.0 17.03.9 4$\times$32$\times$600 38.04.5 0.00.0 9.03.3 12.00.0 0.00.0 3.01.7 36.00.0 5.01.7 21.01.7 16$\times$32$\times$600 44.0 0.0 24.0 12.0 0.0 4.0 36.0 8.0 24.0
@@
When we evaluate more complex compositions using Ineq-Mix, where two seed problems are combined using compositional rules, the difficulty increases further (Figure 2, left). Out of 100 such problems, all whole-proof generation models except DeepSeek-Prover-V2-7B solve fewer than 8, while DeepSeek-Prover-V2-7B significantly leads with 22 solves under a 128-trial budget.
@@
This gap underscores a major limitation: while models may recognize the right approach in natural language, they struggle to implement it in formal proof steps. See Appendix F for examples.
@@
Ineq-Comp offers a complementary lens to existing benchmarks like MiniF2F. While strong models such as DeepSeek-Prover-V2-7B and Kimina-Distill-7B perform well on both, discrepancies emerge that reveal what each benchmark truly measures. For example, Goedel-Prover-SFT (Lin et al., 2025) slightly outperforms STP (without MiniF2F validation augmentation) (Dong and Ma, 2025) on MiniF2F, but is surpassed by STP on Ineq-Comp, which targets compositional generalization. Similarly, InternLM2.5-StepProver+BF (Wu et al., 2024) performs comparably to both on MiniF2F, but struggles significantly on Ineq-Comp, especially on Type I problems, revealing weak high-level compositional reasoning.

### v2

@@ Abstract
LLM-based formal proof assistants (e.g., in Lean) hold great promise for automating mathematical discovery. But beyond syntactic correctness, do these systems truly understand mathematical structure as humans do? We investigate this question in context of mathematical inequalities—specifically the prover’s ability to recognize that the given problem simplifies by applying a known inequality such as AM/GM. Specifically, we are interested in their ability to do this in a compositional setting where multiple inequalities must be applied as part of a solution. We introduce Ineq-Comp, a benchmark built from elementary inequalities through systematic transformations, including variable duplication, algebraic rewriting, and multi-step composition. Although these problems remain easy for humans, we find that most provers—including Goedel, STP, and Kimina-7B—struggle significantly. DeepSeek-Prover-V2-7B shows relative robustness, but still suffers a 20% performance drop (pass@32). Even for DeepSeek-Prover-V2-671B model, the gap between compositional variants and seed problems exists, implying that simply scaling up the model size alone does not fully solve the compositional weakness. Strikingly, performance remains poor for all models even when formal proofs of the constituent parts are provided in context, revealing that the source of weakness is indeed in compositional reasoning. Our results expose a persisting gap between the generalization behavior of current AI provers and human mathematical intuition. All data and evaluation code can be found at https://github.com/haoyuzhao123/LeanIneqComp.
@@ 1 Introduction
Parallel to progress in natural language reasoning, formal theorem proving with LLMs has developed rapidly in the past years. In particular, AlphaProof [^1] https://deepmind.google/discover/blog/ai-solves-imo-problems-at-silver-medal-level/, AlphaGeometry (Trinh et al., 2024), and Seed-Prover (Chen et al., 2025) showed that a model could achieve silver-medal performance in the International Math Olympiad by finding verifiable proofs. Open-source research efforts also show significant progress through both whole-proof generation and tree-search approaches (Xin et al., 2024; Lin et al., 2025a; Dong and Ma, 2025; Wu et al., 2024; Li et al., 2024; Xin et al., 2025; Wang et al., 2025; Lin et al., 2025b), marking rapid advancements in formal theorem proving.
@@
    We develop Ineq-Comp, a benchmark with 75 seed problems solved using basic tools such as AM-GM, Cauchy, and Jensen inequalities and accompanied by a verified Lean 4 proof, and 150 variants generated through compositional transformations (Figure˜1 Left and Section˜2.1). Ineq-Comp is designed to be fully extensible: further composed inequalities can be automatically generated using rule-based transformations on any seed set (Section˜2.2). We also include 50 real-world inequality problems to enhance diversity (Section˜2.3).
@@
    We benchmark a range of LLM-based provers—from general-purpose models like R1-Distill-Qwen to tree-search-based systems—on Ineq-Comp. All models, except DeepSeek-Prover-V2-7B (Ren et al., 2025), suffer from significant performance drops even on the simplest transformations (Section˜3). Notably, DeepSeek-Prover-V2-7B also shows over 20% drop when tested under a limited computational budget (pass@32). These failures persist even when the model is scaled up (Section˜4.1) or provided with the formal proof of the seed problem in-context (Section˜4.2).
@@
    Through detailed output analysis, we identify two key limitations: (1) models rarely attempt to apply core inequalities such as AM-GM or Cauchy within their formal proofs, even when they refer to these techniques in natural language comments; and (2) most models rely heavily on a narrow set of tactics—particularly the sum-of-squares approach using nlinarith—indicating limited diversity in their proof strategies when it comes to inequalities (Section˜3).
@@
    We find that simple fine-tuning on synthetic compositional problems fails to generalize.** While performance improves on problems similar to the training data, the skill does not transfer to new seed problems, suggesting that data augmentation alone is insufficient to teach robust compositional reasoning.
@@ 2 Dataset Curation
    Ineq-Simp (Section˜2.1): 150 problems created by applying human-intuitive transformations to the 75 seed problems. These include algebraic rewrites and compositional constructions (e.g., duplicating variables or multiplying inequalities), as illustrated in Figure˜1 (Left). Problems are categorized by the technique used in the seed (AM-GM, Cauchy, or Miscellaneous).
@@
    Ineq-Mix (Section˜2.2): A rule-based, automated framework that generates new composed inequalities from any given seed set. This enables scalable extension of Ineq-Simp while preserving formal correctness.
@@
    Ineq-Real (Section˜2.3): A set of 50 real-world inequality problems collected from math contests and educational resources to enhance benchmark diversity and assess generalization beyond synthetic construction.
@@ 2.2 Ineq-Mix: A fully expandable benchmark for composition
We introduce Ineq-Mix, an automated framework that generates a potentially unbounded set of valid inequality problems by applying predefined transformation rules to seed problems. These rules fall into three categories: (1) Compositional transformations (e.g., combining inequalities via addition, multiplication, max/min); (2) Variable-level transformations (e.g., replacing variables with algebraic expressions); (3) Problem-level transformations (e.g., applying a monotonic function like $\exp$ or $\log$ to both sides). All generated problems are guaranteed to be mathematically valid and formally solvable. Unlike the simpler variants in Section˜2.1, Ineq-Mix produces more complex compositions that often require nontrivial decomposition and planning. For example, proving $f_{1}f_{2}\geq g_{1}g_{2}$ from $f_{1}\geq g_{1}$ and $f_{2}\geq g_{2}$ may require identifying intermediate inequalities. We generate 100 evaluation problems using only compositional rules (Table˜4), which already challenge state-of-the-art provers. As models improve, Ineq-Mix can scale in complexity by combining multiple transformation types.
To assess whether this difficulty is inherent to formal reasoning, we test OpenAI’s O3 model, which operates in natural language, on problems involving one transformation from each category. O3 reliably decomposes and solves these using intuitive reasoning, with failures typically due to mistakes on individual subproblems. This contrast highlights a key gap: informal models like O3 handle compositionality with ease, while formal provers struggle even on problems built from elementary components. This suggests the bottleneck lies in current formal reasoning systems, not the mathematical content. See Appendix˜B for a full list of transformations used in Ineq-Mix.
@@ 3.1 Experiment setup
We evaluate three categories of models and methods: (1) General-purpose language models, (2) Whole-proof generation models, and (3) Tree-search methods. For general-purpose models, we test DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025) in two settings: with and without explicit thinking steps. Without the chat-style prompt, the model directly completes the Lean 4 proof code. In contrast, when provided with a chat template and generation prompt, the model first performs a “thinking process” in natural language before outputting the full Lean script. For whole-proof generation, we include DeepSeek-Prover-V1.5-RL (Xin et al., 2024), Goedel-Prover-SFT (Lin et al., 2025a), STP (Dong and Ma, 2025), Kimina-7B (Wang et al., 2025), and the latest state-of-the-art model DeepSeek-Prover-V2-7B (Ren et al., 2025), which achieves top performance on the MiniF2F benchmark. For tree-search-based approaches, we evaluate DeepSeek-Prover-V1.5-RL+RMaxTS (Xin et al., 2024) and InternLM-2.5-Step-Prover+BF (Wu et al., 2024). While recent models such as HunyuanProver (Li et al., 2024) and BFS-Prover (Xin et al., 2025) show stronger performance on MiniF2F, we exclude them due to the lack of publicly available code.
@@
We evaluate all methods using the standard pass@$N$ accuracy, where a problem is considered solved if at least one of $N$ generated proofs is correct. For tree-search methods, budget denotes the total number of interactions with the Lean compiler. For additional details, see Appendix˜D, and refer to Appendix˜E for the prompt templates used in evaluation.
@@ 3.2 Main findings and discussion
The experiment results on Ineq-Simp: categorized by the natural of seed problems Ineq-AMGM, Ineq-Cauchy, and Ineq-MISC are summarized in Table˜1. The performance of different models on Ineq-Mix (randomly generated) and Ineq-Real is shown in Figure˜2.
@@
Our experiments reveal a clear failure of compositional generalization across most current theorem provers, even on problems derived from simple seed inequalities that remain easy for humans. As shown in Table 1, top models such as Goedel-Prover-SFT (Lin et al., 2025a), STP (Dong and Ma, 2025), and Kimina-Prover (Wang et al., 2025) show large drops in accuracy from seed problems to their compositional variants: Type I (duplication-based) and Type II (algebraic transformation).
For instance, Kimina-Prover-Preview-Distill-7B achieves 80% on AM-GM seed problems (budget 3200), but drops to 44% on Type I and 64% on Type II. The drop is even more dramatic for Goedel-Prover and DeepSeek-Prover-V1.5, which struggle to exceed 5% accuracy on Ineq-AMGM Type I problems—even with a large number of attempts. The only partial exception is DeepSeek-Prover-V2-7B, which reduces the seed-to-Type I gap to 13% and seed-to-Type II to 18%. Its drop becomes more pronounced at smaller budgets (e.g., pass@32), confirming that the challenge is not just probabilistic but fundamentally hard (Figure˜1, right). This is especially striking given that DeepSeek-Prover-V2 is explicitly trained using a divide-and-conquer strategy (Ren et al., 2025), which should, in principle, help with decomposing Type I problems. Its performance gap, particularly under low-pass settings, highlights the intrinsic difficulty of compositional reasoning in formal reasoning.
@@
[…]
@@
When we evaluate more complex compositions using Ineq-Mix, where two seed problems are combined using compositional rules, the difficulty increases further (Figure˜2, left). Out of 100 such problems, all whole-proof generation models except DeepSeek-Prover-V2-7B solve fewer than 8, while DeepSeek-Prover-V2-7B significantly leads with 22 solves under a 128-trial budget.
@@
This gap underscores a major limitation: while models may recognize the right approach in natural language, they struggle to implement it in formal proof steps. See Appendix˜F for examples.
@@
Ineq-Comp offers a complementary lens to existing benchmarks like MiniF2F. While strong models such as DeepSeek-Prover-V2-7B and Kimina-Distill-7B perform well on both, discrepancies emerge that reveal what each benchmark truly measures. For example, Goedel-Prover-SFT (Lin et al., 2025a) slightly outperforms STP (without MiniF2F validation augmentation) (Dong and Ma, 2025) on MiniF2F, but is surpassed by STP on Ineq-Comp, which targets compositional generalization. Similarly, InternLM2.5-StepProver+BF (Wu et al., 2024) performs comparably to both on MiniF2F, but struggles significantly on Ineq-Comp, especially on Type I problems, revealing weak high-level compositional reasoning.